Problem 9-5 Option to Walt Your company is deciding whether to invest in a new machine. The new machine wil increase cash flow by $326,000 per year. You believe the technology used in the machine has a 10-year life; in other words, no matter when you purchase the machine, it will be obsolete 10 years from today. The machine is currently priced at $1,760,000. The cost of the machine will decline by $111,000 per year until it reaches $1,205,000, where it will remain If your required return is 13 percent, calculate the NPV if you purchase the machine today. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16) NPV$ If your required return is 13 percent, calculate the NPV if you wait to purchase the machine until the indicated year. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) NPV Year 1 Year 2 Year 3 Year 4 Year 5 Year 6 Should you purchase the machine? OYes ONo If so, when should you purchase it? O Today OOne year from now OTwo years from now
Answer #1

Solution:

i. If we purchase the machine today, the NPV is

NPV = -1,760,000 + 326,000 (PVIFA @ i, n)

NPV = -1,760,000 + 326,000 (PVIFA @ 13%, 10)

NPV = -1,760,000 + 326,000 [(1.13^10-1)/(0.13*1.13^10)]

NPV = -1,760,000 + 326,000 (5.4262)

NPV = $8,955.37

ii. Year 1

NPV = [-1,649,000 + 326,000 (PVIFA @ 13%, 9)]/1.13

NPV = [-1,649,000 + 326,000 (5.1317)]/1.13

NPV = $21,167.76

Year 2

NPV = [-$1,538,000 + 326,000 (PVIFA @ 13%, 8)]/1.13^2

NPV = [-$1,538,000 + 326,000 (4.7988)]/1.13^2

NPV = $20,674.38

Year 3

NPV = [-$1,427,000 + 326,000 (PVIFA @ 13%, 7)]/1.13^3

NPV = [-$1,427,000 + 326,000 (4.4226)]/1.13^3

NPV = $10,237.04

Year 4

NPV = [-$1,316,000 + 326,000 (PVIFA @ 13%, 6)]/1.13^4

NPV = [-$1,316,000 + 326,000 (3.9978)]/1.13^4

NPV = -$7,849.72

Year 5

NPV = [-$1,205,000 + 326,000 (PVIFA @ 13%, 5)]/1.13^5

NPV = [-$1,205,000 + 326,000 (6.4082)]/1.13^5

NPV = -$31,687.74

Year 6

NPV = [-$1,205,000 + 326,000 (PVIFA @ 13%, 4)]/1.13^6

NPV = [-$1,205,000 + 326,000 (2.9745)]/1.13^6

NPV = -$113,029.68

iii. Yes, you should purchase the machine because NPV is positive.

iv. Machine should be purchased One Year from Now because NPV is highest.